On biplanes (56,11,2) with automorphism groups of order four

Vladimir Ćepulić, Zagreb, Croatia

Keywords: biplane, automorphism group, orbutal structure, indexing AMS Subject classification: 05B05

Abstract: We prove that there is no biplane (56,11,2) admitting an automorphism group of order 4 operating semiregularly. This completes the classification of such biplanes admitting an automorphism of order 4.